# Isometry groups of Polish ultrametric spaces

### v2

@@ 1. Introduction
An ultrametric $d$ on a set $X$ is a metric satisfying the following strong form of the triangle inequality, for all $x,y,z\in X$:
@@
In the 1950’s, the algebraic number theorist Marc Krasner posed the following problem [Kra56, LS86]: [^1] A related problem posed by Pestov, asking for a characterization of all subgroups of the isometry groups of ultrametric spaces, was solved by Lemin and Smirnov [LS86] in 1986.
@@
This is one of the oldest problems concerning ultrametric spaces, as it was proposed shortly after the concept of ultrametric was introduced by Krasner himself as an abstraction of the special properties of the $p$-adic metrics [Kra44]. Although the problem has remained wide open for about 70 years, several partial results were obtained, each one involving some form of generalized wreath product of transitive permutation groups. For example, Feĭnberg [Fei74, Fei78] showed that if $X$ is a compact ultrametric space, then its isometry group is isomorphic to a (unrestricted) generalized wreath product, denoted by $\operatorname{Wr}^{\operatorname{Max}}_{\delta\in\Delta}H_{\delta}$ in this work, as defined by Holland [Hol69]. This result was later extended by Nosova and Feĭnberg [NF80] to all $T$-complete (also called spherically complete) ultrametric spaces. More recently, Malicki considered what he calls $W$-spaces, i.e. Polish ultrametric spaces which are locally non-rigid and exact, [^2] See Section 6.2 for the definition. and showed that their isometry groups can again be characterized in terms of generalized wreath products, this time using a suitable variant of Holland’s construction that in this paper is denoted by $\operatorname{Wr}^{\operatorname{Wsp}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ [Mal14]. We remark that Krasner’s problem and all the partial results mentioned so far are purely algebraic, as no attempt to define a group topology on generalized wreath products appears in the literature.
Problem 1 reappeared in 2003 in a slightly different context, namely, that of topological groups. The isometry group $\operatorname{Iso}(X)$ of any Polish (i.e. separable complete) metric space is a Polish group when equipped with the pointwise convergence topology. In their seminal paper [GK03], Gao and Kechris proved that, indeed, every Polish group is of this form, up to topological group isomorphism; therefore, the class of isometry groups of Polish metric spaces and the class of Polish groups are the same. This naturally started the search for similar characterizations for interesting subclasses of Polish metric spaces. To mention a few examples: compact Polish groups are the isometry groups of compact metric spaces [GK03, Mel08]; closed subgroups of $\operatorname{Sym}(\omega)$ are the isometry groups of zero-dimensional locally compact Polish metric spaces [GK03]; locally compact Polish groups are the isometry groups of proper Polish metric spaces [GK03, MS09].
@@
The main goal of this paper is to solve Krasner’s Problem 1 for Polish ultrametric spaces and, simultaneously, to fully solve Gao-Kechris’ Problem 2 for both arbitrary Polish ultrametric spaces and for the locally compact ones — indeed, we will see that there is no difference between these two variants of the problem. The main ingredients we use are the following.
@@
In Section 5 we move instead to (unrestricted) generalized wreath products. First we present the variants already considered in the literature, namely, Hall’s wreath product $\operatorname{Wr}^{\operatorname{Fin}}_{\delta\in\Delta}H_{\delta}$ from [Hal62], Holland’s wreath product $\operatorname{Wr}^{\operatorname{Max}}_{\delta\in\Delta}H_{\delta}$ from [Hol69], and Malicki’s wreath product $\operatorname{Wr}^{\operatorname{Wsp}}_{\delta\in\Delta}H_{\delta}$ from [Mal14], in the unified framework of generalized wreath products over a global domain $\operatorname{Wr}^{S}_{\delta\in\Delta}H_{\delta}$ (Section 5.1). Then we consider a fourth option based on locally finite supports $\operatorname{Wr}^{\operatorname{LF}}_{\delta\in\Delta}H_{\delta}$, and show that, among the mentioned four variants, this is the only class giving rise to Polish groups when we equip the generalized wreath products with a very natural topology (see Proposition 5.5 and the discussion following it). Finally, we introduce two powerful generalizations of the above mentioned generalized wreath products, that is, generalized wreath products over local domains $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}H_{\delta}$ (Section 5.2) and projective wreath products $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}H_{\delta}$ (Section 5.3).
Combining these tools all together, in Section 6 we obtain the desired characterization of isometry groups of (locally compact) Polish ultrametric spaces in terms of generalized wreath products of full permutation groups.
@@
Our methods are flexible enough to allow for restricted forms of the above characterization, drawing exact correspondences between certain natural subclasses of Polish ultrametric spaces (like homogeneous or homogeneous discrete spaces, Urysohn spaces, and exact spaces), and various natural variants (both old and new) of generalized wreath products. The main results are summarized in the table from Appendix A. For example, all characterizations can be phrased using only the variant introduced in this paper, namely, the one based on locally finite supports. Or we can determine to what extent the generalized wreath products previously appeared in the literature can be used in this context:
@@
Along the way, we also provide results which might be of independent interest in the context of group theory. For example, we show that, under mild conditions, projective wreath products and generalized wreath products over local domains give rise to the same class of groups (see Theorems 5.12 and 6.15), and we identify natural conditions under which generalized wreath products over local domains can be realized as the more classical generalized wreath products over global domains (Proposition 5.9). We conclude the paper with some additional results and a few suggestions for future work (Section 7).
@@ 3. $L$-trees
Notice that if the reverse of $L$ is well-founded (or even just $\omega+1$ does not embed into $L$), then every weak $L$-tree is an $L$-tree.
@@ 7. Additional results and open problems
The characterizations obtained in this paper might be helpful in tackling some natural open problems in the area. For example, Gao and Kechris showed that if the Polish ultrametric space $X$ is Heine-Borel, then $\operatorname{Iso}(X)$ is the closure of an increasing union of compact subgroups, and hence it is amenable ([GK03, Theorem 8.9]). This raises the question of which other isometry groups of Polish ultrametric spaces are amenable. Similar questions might be asked for interesting group properties other than amenability, such as being oligomorphic, or having uncountable strong cofinality. Concerning the latter, for example, one might attempt to generalize Malicki’s characterization [Mal14, Theorem 5.3], which crucially uses generalized wreath products, from $W$-spaces to arbitrary Polish ultrametric spaces.
@@ Appendix A Summarizing table
[…]
$\mathbf{S}$ family of local domains: $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}$ $\big{[}\mathbf{S}\subseteq\mathbf{S}^{\operatorname{Max}}\big{]}$ $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $\mathbf{S}$ local. homog. (∗) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $\mathbf{S}$ local. homog. (∗) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\mathbf{S}$ full and locally homogeneous (†) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$
[…]
Projective wreath products: $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}$ $\big{[}\mathbf{S}\subseteq\mathbf{S}^{\operatorname{Max}}\big{]}$ $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ loc. homog. (∗) $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ loc. homog. (∗) $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ full and locally homogeneous (†)
$\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$
[…]

### v3

@@ 1. Introduction
An ultrametric (or non-Archimedean metric) $d$ on a set $X$ is a metric satisfying the following strong form of the triangle inequality, for all $x,y,z\in X$:
@@
Ultrametric spaces have grown in importance in mathematics, where, besides descriptive set theory and general topology, they are used e.g. in number theory ($p$-adic analysis) and combinatorics (consider e.g. the distance given by the weight of the minimax path between vertices in an edge-weighted undirected graph). Ultrametrics are widely considered in computer science too, where distances between words (such as the $p$-closed distance) are often ultrametrics, but also in other fields such as condensed matter physics [MPV87], the theory of aperiodic solids [RTV86], the study of certain discrete models of turbulence of fluids [BBT97] or of disordered systems in equilibrium statistical mechanics [PRT00, Par23], data science [Mur07], taxonomy [LL12], and so on.
In the 1950’s, the algebraic number theorist Marc Krasner posed the following problem [Kra56, LS86]: [^1] Surprisingly, a related problem posed by Pestov, asking for a characterization of all subgroups of the isometry groups of ultrametric spaces, turned out to be much simpler than Problem 1, and was solved by Lemin and Smirnov [LS86] already in 1986.
@@
This is one of the oldest problems concerning ultrametric spaces, as it was proposed shortly after the concept of ultrametric was introduced by Krasner himself as an abstraction of the special properties of the $p$-adic metrics [Kra44]. Although the problem has remained wide open for about 70 years, several partial results were obtained, each one involving some form of generalized wreath product of transitive permutation groups. For example, Feĭnberg [Fei74, Fei78] showed that if $X$ is a compact ultrametric space, then its isometry group is isomorphic to a (unrestricted) generalized wreath product, denoted by $\operatorname{Wr}^{\operatorname{Max}}_{\delta\in\Delta}H_{\delta}$ in this work, as defined by Holland [Hol69]. This result was later extended by Nosova and Feĭnberg [NF80] to all $T$-complete (also called spherically complete) ultrametric spaces. More recently, Malicki considered what he calls $W$-spaces, i.e. Polish ultrametric spaces which are locally non-rigid and exact, [^2] See Section 6.2 for the definition. and showed that their isometry groups can again be characterized in terms of generalized wreath products, this time using a suitable variant of Holland’s construction that in this paper is denoted by $\operatorname{Wr}^{\operatorname{Wsp}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ [Mal14]. Thus, Krasner’s Problem 1, which has a somewhat vague formulation, has always been intended as the request of isolating the right notion of generalized wreath product giving rise to the isometry groups of (Polish) ultrametric spaces. We also remark that Krasner’s problem and all the partial results mentioned so far are purely algebraic, as no attempt to define a group topology on generalized wreath products appears in the literature.
Problem 1 reappeared in 2003 in a slightly different context, namely, that of topological groups. The isometry group $\operatorname{Iso}(X)$ of any Polish (i.e. separable complete) metric space is a Polish group when equipped with the pointwise convergence topology. In their seminal paper [GK03], Gao and Kechris proved that, indeed, every Polish group is of this form, up to topological group isomorphism; therefore, the class of isometry groups of Polish metric spaces and the class of Polish groups are the same. This naturally started the search for similar characterizations for interesting subclasses of Polish metric spaces. To mention a few examples: compact Polish groups are the isometry groups of compact metric spaces [GK03, Mel08]; closed subgroups of $\operatorname{Sym}(\omega)$ are the isometry groups of zero-dimensional locally compact Polish metric spaces [GK03]; locally compact Polish groups are the isometry groups of proper Polish metric spaces [GK03, MS09]. Some characterizations are expressed in terms of a specific group operator applied to a certain class of Polish groups: for example, the closed subgroups of groups of the form $\prod_{n\in\omega}(\operatorname{Sym}(\omega)\ltimes G_{n}^{\omega})$, for $(G_{n})_{n\in\omega}$ a sequence of locally compact Polish groups, characterize the isometry groups of locally compact (or, equivalently, $\sigma$-compact) Polish metric spaces [GK03].
@@
The main goal of this paper is to solve Krasner’s Problem 1 for Polish ultrametric spaces (in the algebraic sense, as originally formulated, but also in the finer context of topological groups) and, simultaneously, to fully solve Gao-Kechris’ Problem 2 for both arbitrary Polish ultrametric spaces and for the locally compact ones — indeed, we will see that there is no difference between these two variants of the problem. The main ingredients we use are the following.
@@
In Section 5 we move instead to (unrestricted) generalized wreath products. Although generalized wreath products have been involved in each of the partial results towards the solution of Problem 1 previously mentioned, the variants of this algebraic construction that can be found in the literature fail to provide a complete answer. To overcome this, we introduce a very general approach to wreath products that we believe could pique the interest also of a group theory oriented reader. More in detail, we first present the known variants (namely, Hall’s wreath product $\operatorname{Wr}^{\operatorname{Fin}}_{\delta\in\Delta}H_{\delta}$ from [Hal62], Holland’s wreath product $\operatorname{Wr}^{\operatorname{Max}}_{\delta\in\Delta}H_{\delta}$ from [Hol69], and Malicki’s wreath product $\operatorname{Wr}^{\operatorname{Wsp}}_{\delta\in\Delta}H_{\delta}$ from [Mal14]), in the unified framework of generalized wreath products over a global domain $\operatorname{Wr}^{S}_{\delta\in\Delta}H_{\delta}$ (Section 5.1). Then we consider a fourth option based on locally finite supports $\operatorname{Wr}^{\operatorname{LF}}_{\delta\in\Delta}H_{\delta}$, and show that, among the mentioned four variants, this is the only class giving rise to Polish groups when we equip the generalized wreath products with a very natural topology (see Proposition 5.5 and the discussion following it). Finally, we introduce two powerful generalizations of the above mentioned generalized wreath products, that is, generalized wreath products over local domains $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}H_{\delta}$ (Section 5.2) and projective wreath products $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}H_{\delta}$ (Section 5.3).
Combining these tools all together, in Section 6 we obtain the desired characterization of isometry groups of (locally compact) Polish ultrametric spaces in terms of generalized wreath products of full permutation groups. We remark that the following result can be read in purely algebraic terms (i.e. omitting any topological consideration on the groups and on their isomorphisms), or in the finer setup of Polish groups.
@@
Our characterization is naturally tied to quite old and well-established group theoretic constructions, like generalized wreath products applied to full permutation groups, and it aligns with all previous partial results on the matter in the literature; therefore, we believe that Theorem 1.1 provides an adequate and satisfactory solution to both long-standing Problems 1 and 2 at once. Moreover, our methods are flexible enough to allow for restricted forms of such characterization, drawing exact correspondences between certain natural subclasses of Polish ultrametric spaces (like homogeneous or homogeneous discrete spaces, Urysohn spaces, and exact spaces), and various natural variants (both old and new) of generalized wreath products. The main results are summarized in the table from Appendix A. For example, all characterizations can be phrased using only the variant introduced in this paper, namely, the one based on locally finite supports. Or we can determine to what extent the generalized wreath products previously appeared in the literature can be used in this context:
@@
Along the way, we also provide results that are of independent interest in the context of group theory. For example, we show that, under mild conditions, projective wreath products and generalized wreath products over local domains give rise to the same class of groups (see Theorems 5.12 and 6.15), and we identify natural conditions under which generalized wreath products over local domains can be realized as the more classical generalized wreath products over global domains (Proposition 5.9).
Our neat presentation of isometry groups of Polish ultrametric spaces as generalized wreath products unveil their rich combinatorial structure and, together with the concepts and methods developed to study them, may unlock the solution to various key problems in the area concerning the properties of such groups and their actions. We collect some of these problems, together with a few additional results, in Section 7.
@@ 3. $L$-trees
Notice that if the reverse of $L$ is well-founded (or even just $\omega+1$ does not embed into $L$), then every weak $L$-tree is an $L$-tree. Picture 1 depicts an $L$-tree, together with the main ingredients involved in its definition.
[…]
Figure 1. A graphical representation of an arbitrary $L$-tree, highlighting in particular the splitting level $\mathrm{split}(t,t^{\prime})$ of the two $\leq_{T}$-incomparable nodes $t$ and $t^{\prime}$.
@@ 7. Additional results and open problems
The characterizations presented in this paper might provide a new tool to understand the complexity under Borel reducibility of the isomorphism relation on classes of non-Archimedean Polish groups (that is, closed subgroups of $\operatorname{Sym}(\omega)$), a program that started by Kechris, Nies and Tent in [KNT18]. In particular, one problem considered in this line of research is establishing the complexity of the isomorphism relation on isometry groups of Polish ultrametric spaces. As a test case, one could first restrict the attention to Polish ultrametric Urysohn spaces. A related problem was already considered in [EGLMM21, EGL24], where the authors studied the isomorphism types of isometry groups of $\Delta$-metric Urysohn spaces, and showed that they are completely determined by their distance set $\Delta$ (up to a natural equivalence notion). It is natural to ask if a similar phenomenon arises also in the context of isometry groups of Polish ultrametric Urysohn spaces. Our structural results, and in particular Theorem 6.7, may help to show that there are many non-isomorphic such groups, as a step to prove that the isomorphism relation on them is very complicated.
Our results might also be helpful in tackling some natural open problems in the context of topological groups. For example, Gao and Kechris showed that if the Polish ultrametric space $X$ is Heine-Borel, then $\operatorname{Iso}(X)$ is the closure of an increasing union of compact subgroups, and hence it is amenable ([GK03, Theorem 8.9]). This raises the question of which other isometry groups of Polish ultrametric spaces are amenable. Similar questions might be asked for interesting group properties other than amenability, such as being oligomorphic, or having uncountable strong cofinality. Concerning the latter, for example, one might attempt to generalize Malicki’s characterization [Mal14, Theorem 5.3], which crucially uses generalized wreath products, from $W$-spaces to arbitrary Polish ultrametric spaces.
@@ Appendix A Summarizing table
$\operatorname{Iso}(U)$ for $U$ Polish ultrametric space $U$ homogeneous and unif. discrete $U$ homogeneous $U$ exact all $U$
$U$ exact and unif. discrete $U$ unif. discrete
$U$ exact and perfect loc. compact $U$ perfect loc. compact
$\operatorname{Aut}(\mathcal{T})$ for $\mathcal{T}$ countable pruned $L$-tree $\mathcal{T}$ homogeneous and $L$ has $\min$ $\mathcal{T}$ homogeneous $\mathcal{T}$ $L$-tree with ($\star$)6.2 all $\mathcal{T}$
$\mathcal{T}$ special $L$-tree with ($\star$)6.2 and $L$ has $\min$ $\mathcal{T}$ special $L$-tree and $L$ has $\min$
$\mathcal{T}$ special $\mathbb{Q}$-tree with ($\star$)6.2 $\mathcal{T}$ special $\mathbb{Q}$-tree
Locally finite supports: $\operatorname{Wr}^{\operatorname{LF},-}_{\delta\in\Delta}$ $\operatorname{Wr}^{\operatorname{LF}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $\operatorname{Wr}^{\operatorname{LF}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $\operatorname{Wr}^{\operatorname{LF}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ (†) $\operatorname{Wr}^{\operatorname{LF},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ (‡)
“Classical” generalized wreath products: $\operatorname{Wr}^{\operatorname{Fin}/\operatorname{Wsp}}_{\delta\in\Delta}$ Hall’s groups with topology $\tau^{*}$ Malicki’s groups with $\Delta$ linear Malicki’s groups (†)
Hall’s groups $\Delta$ with $\min$
Malicki’s groups $\Delta$ linear with $\min$
$S$ global domain: $\operatorname{Wr}^{S}_{\delta\in\Delta}$ $\big[S\subseteq S^{\operatorname{Max}}\big]$ $\operatorname{Wr}^{S}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $S$ appr. homog. $\operatorname{Wr}^{S}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $S$ appr. homog. $\operatorname{Wr}^{S}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $S$ appr. homog. (†)
$\mathbf{S}$ family of local domains: $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}$ $\big[\mathbf{S}\subseteq\mathbf{S}^{\operatorname{Max}}\big]$ $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $\mathbf{S}$ local. homog. (∗) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $\mathbf{S}$ local. homog. (∗) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\mathbf{S}$ full and locally homogeneous (†) $\operatorname{Wr}^{\mathbf{S}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ (♯)
Projective wreath products: $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}$ $\big[\mathbf{S}\subseteq\mathbf{S}^{\operatorname{Max}}\big]$ $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear with $\min$ $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ loc. homog. (∗) $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\Delta$ linear $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ loc. homog. (∗) $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ $\langle\mathbf{S},\boldsymbol{\pi}\rangle$ full and locally homogeneous (†) $\operatorname{Wr}^{\mathbf{S},\boldsymbol{\pi}}_{\delta\in\Delta}\operatorname{Sym}(N_{\delta})$ (♯)